# Complex Domain Approach for Reversible Data Hiding and Homomorphic Encryption: General Framework and Application to Dispersed Data

### v2

@@ 1 Introduction
The ubiquitous generation and pervasive use of digital data stand as defining characteristics of the modern era. From vast networks of Internet of Things (IoT) devices and industrial sensors to intricate smart city infrastructures and critical healthcare systems, data fuels decision-making, automates processes, and enables unprecedented levels of connectivity. This reliance on data, often collected from distributed and potentially vulnerable environments, is critically linked to its trustworthiness (Yan et al., 2014). Ensuring the confidentiality, integrity, authenticity, and privacy of information has become a priority, as compromised data can lead to erroneous conclusions, system failures, and severe breaches of privacy or security (Liang et al., 2016).
@@
Within this complex digital landscape, the concept of Reversible Data Hiding (RDH) (Shi et al., 2016; Kumar et al., 2022) emerges as a specialized and increasingly popular approach to data security. RDH involves embedding secret information (a “watermark”) directly into a host data item in such a way that, after extraction of the watermark, the original host data can be perfectly and exactly restored to its original state. This unique property, combining data embedding with guaranteed reversibility, makes RDH appropriate in sensitive applications such as medical imaging, legal documentation, and military intelligence, where any permanent alteration of the original data is unacceptable. Unlike traditional data encryption, which primarily ensures confidentiality, RDH focuses on providing intrinsic integrity, authenticity, and provenance, within the data itself.
@@
In the context of individual, scalar, or time-series numerical data, such as those originating from Wireless Sensor Networks (WSNs) (Shi and Xiao, 2013), IoT devices (Rahman et al., 2021), or critical financial and medical monitoring systems (Bhalerao et al., 2019), RDH offers unique advantages. Foremost among these is its ability to ensure robust data integrity and authenticity at the granular level of single readings or sequential observations. By embedding a watermark that can be extracted and verified, receivers (e.g., a Data Collector in a WSN) can confirm that the data originated from a legitimate source and has not been tampered with in transit. This inherent capability to facilitate data provenance and traceability is particularly valuable, allowing for forensic analysis of data origin and processing paths in distributed and often resource-constrained environments.
@@
Beyond these general benefits, RDH offers distinct advantages compared to other commonly employed security primitives in certain contexts. Unlike cryptographic hashing or digital signatures, which typically append a separate authentication tag or signature to the data, RDH embeds the integrity verification mechanism directly within the data itself. This can simplify data handling and ensure the integrity proof travels jointly with the content. While conventional encryption ensures confidentiality, it does not directly provide data integrity; and Physical Unclonable Functions (PUFs) (Gao et al., 2020) focus on device authentication rather than data integrity. RDH uniquely combines content-dependent integrity (through watermark verification) with the guarantee of perfect original data restoration, a feature not offered by one-way hash functions or irreversible authentication tags. This intrinsic embedding can be particularly beneficial in constrained environments or specialized data formats where appending additional metadata is undesirable or challenging.
@@
In a parallel development within cryptography, Homomorphic Encryption (HE) (Acar et al., 2018) has emerged as a transformative technology for data privacy (Iezzi, 2020). HE schemes allow computations to be performed directly on encrypted data without ever needing to decrypt it (Zhang et al., 2020). The output of such a computation is an encrypted result that, when decrypted, is identical to the result of the same computation performed on the unencrypted data. This property enables the processing of sensitive information by untrusted third parties (such as cloud servers) while ensuring that the data’s confidentiality remains fully protected throughout its lifecycle.
@@ 1.1 Plan and Contributions of the Paper
  • A detailed novel extension of ElGamal encryption to the Gaussian integers, serving as a foundational cryptographic primitive for the complex domain.
  • The H[i][i]dden-EG protocol, providing joint reversible information hiding and encryption specifically for individual data points, making use of the extended Gaussian ElGamal scheme.
@@ 2.1 Reversible Data Hiding
Reversible Data Hiding (Shi et al., 2016; Kumar et al., 2022) is a specialized field within information hiding that focuses on embedding secret information into host data in such a way that the original host data can be perfectly and exactly recovered after the hidden data (watermark) has been extracted. This property is essential in applications that do not allow the permanent alteration of the original content.
Historically, RDH techniques have been extensively developed and applied primarily to multimedia content, including digital images, audio, and video (Shi et al., 2016). Methods like Least Significant Bit (LSB) modification (when combined with original LSB storage) (Celik et al., 2002), Difference Expansion (DE) (Tian, 2003), Histogram Shifting (HS) (Ni et al., 2006), and Prediction-Error Expansion (PEE) (Ou et al., 2013) use the inherent statistical redundancy and the human perceptual system’s tolerance for slight alterations in these rich data formats. These approaches can embed information in an imperceptible mmanner while ensuring perfect reversibility, meaning the host data is faithfully restored to its original state after hidden data extraction.
@@
For this class of individual or dispersed data, existing reversible embedding methods tend to fall into a few categories. One common approach is Modulo or Remainder-based embedding, where the secret information $w$ (an integer in range $0,\ldots,q-1$) is embedded into a host integer $d$ using an operation like $d^{\prime}=q\cdot d+w$ (Kabir et al., 2025). A common special case involves embedding a single bit ($w\in\{0,1\}$) by setting $q=2$, resulting in $d^{\prime}=2d+w$ (Kabir et al., 2024). Upon extraction, $w$ is simply $d^{\prime}\,(\operatorname{\mathrm{mod}}q)$ and $d$ is $\lfloor d^{\prime}/q\rfloor$. When $q$ is a power of 2 ($q=2^{B}$), these extraction operations can be efficiently performed using bitwise logic (e.g., $w=d^{\prime}\text{ \& }(q-1)$ and $d=d^{\prime}\gg B$).
@@
Another relevant approach for sequential data is Difference Expansion (DE), or its more general form, Prediction-Error Expansion (Shi and Xiao, 2013). These methods exploit relationships between consecutive data samples by computing prediction errors (e.g., $e_{k}=d_{k}-d_{k-1}$) and then embedding information by expanding these errors.
@@ 2.2 Homomorphic Encryption
Homomorphic Encryption (Acar et al., 2018) is a unique and powerful class of cryptographic schemes that allows a user to perform computations on encrypted data without first having to decrypt it. The result of these operations is an encrypted output which, when decrypted, is the same as if the operations had been performed on the original, unencrypted data. This remarkable capability is useful for maintaining data privacy and confidentiality when computations are outsourced to untrusted third parties, such as cloud services or remote data aggregators.
HE schemes are typically classified based on the types and number of operations they support. Partially Homomorphic Encryption (PHE) schemes, such as (Cominetti and Simplicio, 2020), support an unlimited number of a single type of operation, either addition or multiplication. Somewhat Homomorphic Encryption (SHE) schemes, which include the pioneering work by Boneh-Goh-Nissim (BGN) in (Boneh et al., 2005), can perform a limited number of both additions and multiplications. Finally, Fully Homomorphic Encryption (FHE) schemes, such as Ring Learning with Errors (RLWE) (Brakerski and Vaikuntanathan, 2011), support an unlimited number of both operations, allowing for arbitrary computations on encrypted data. In this paper, two classical PHE schemes, ElGamal (1985) and Paillier (1999), are applied for their distinct homomorphic properties that enable the proposed protocols for both individual data encryption and privacy-preserving data aggregation.
@@ 2.2.1 ElGamal Encryption
ElGamal encryption (ElGamal, 1985) is an asymmetric key encryption algorithm based on the Diffie-Hellman key exchange scheme. Its security relies on the difficulty of the discrete logarithm problem (DLP) in the multiplicative group of integers modulo a prime number.
@@ 2.2.2 Paillier Encryption
Paillier (1999) is a probabilistic asymmetric encryption algorithm. Unlike ElGamal, which is homomorphic with respect to multiplication, Paillier encryption is additively homomorphic, meaning that sums of plaintexts can be computed from their ciphertexts.
@@ 3.1 Basic Definitions
Unlike reversible data hiding for multimedia content or highly redundant data —where source redundancy allows for techniques such as difference expansion, histogram shifting, or prediction error expansion to hide information without necessarily increasing carrier size— embedding a watermark $w_{k}$ in a single scalar $d_{j,k}$ such that both can be perfectly recovered requires data expansion. For instance, as seen in Kabir et al. (2025), embedding a $B$-bit watermark using an operation like $d^{\prime}_{j,k}=2^{B}\cdot d_{j,k}+w_{k}$ increases the bit-length of the original data by exactly $B$ bits. From an information-theoretic perspective, if the host data consists of $A$ bits and the watermark comprises $B$ bits, and both are independent and occupy their full possible ranges, the combined state space consists of $2^{A+B}$ distinct values. In such cases, the result must be encoded with at least $A+B$ bits for perfect recovery. While “slack” in a data’s restricted range (e.g., values $d_{j,k}\in[101,200]$ encoded in a 1-byte container) can theoretically host bits without expansion, scalar reversible embedding of independent, full-range variables fundamentally requires increasing the bit-length of the transmitted values. $\triangle$
@@ 3.2 Proposed Reversible Watermarking Scheme
The proposed method is vulnerable against Gaussian integer factorization. Even with just one eavesdropped value, a malicious observer may try to compute the Gaussian integer factors of the watermarked data and obtain knowledge about the original data, the watermark or the watermarking key. Consequently, for practical applications, it is not recommended to rely solely on the challenge $\lambda_{k}$ as a security measure. The scheme can be combined with random numbers and/or cryptographic functions for increased security and privacy, as shown in the protocols proposed in Section 5.
$\triangle$
@@
Hence, an eavesdropper who accessed $\delta^{\prime}_{j,k}$ would not be able to determine the data or the watermark directly, but may try to infer information from Gaussian integer factorization, as discussed in Remark 8. In addition, if the eavesdropper had access to the watermarked values corresponding to different time samples ($k$) or different devices ($j$), they could try to obtain even more information about $\lambda_{k}$, especially if it is not changed frequently.
$\triangle$
@@ 4 Homomorphic Encryption on the Complex Domain
The extension of ElGamal encryption to Gaussian integers requires a careful examination of the mathematical basis of the cryptosystem. The first building block of the standard ElGamal cryptosystem is the multiplicative cyclic group $\mathbb{Z}^{*}_{p}$, for a prime number $p$, and a generator $g$ for such a group. In the Gaussian integer case, we must have a prime number $\pi$ in $\mathbb{Z}[i]$ and a consistent definition of the modular arithmetic $(\operatorname{\mathrm{mod}}\pi)$.
For a better understanding of the adaptation of this cryptosystem to the complex domain, a brief introduction to modular arithmetic over Gaussian integers is provided in A. A more detailed and comprehensive reference on Gaussian integers can be found in (Conrad, 2016).
@@ 4.1 ElGamal Encryption over Gaussian Integers
The principles of ElGamal encryption can be extended to other finite cyclic groups, including the multiplicative group of Gaussian integers modulo a Gaussian prime. This adaptation exploits the rich algebraic structure of Gaussian integers while maintaining the security reliance on the DLP.
@@
As a key property, the order of this group, $\mathcal{N}=p^{2}-1$, is significantly larger than p1p-1 for the same numerical prime $p$ used in the regular integer ElGamal. This relationship, $\mathcal{N}=(p-1)(p+1)$, is particularly advantageous for security. If a prime $p$ is chosen such that p1p-1 possesses a large prime factor (a standard requirement for robust security in integer-based ElGamal against algorithms like Pohlig-Hellman (Pohlig and Hellman, 1978)), then this same large prime factor will inherently be a factor of $\mathcal{N}$. Therefore, the Gaussian integer ElGamal variant naturally inherits at least the same level of security (in terms of resistance to Pohlig-Hellman attacks) as its regular integer counterpart for a given $p$. Furthermore, the existence of the additional factor $(p+1)$ in $\mathcal{N}$ can potentially introduce even larger prime factors, further enhancing the security of the scheme. Thus, careful selection of the Gaussian prime $p$ to ensure that the largest prime factor of $\mathcal{N}$ is sufficiently large guides robust parameter choices.
@@ 7 Conclusion
This paper introduces Hiding in the Imaginary Domain with Data Encryption (H[i][i]dden), a novel framework for RDH that operates uniquely within the complex domain. By intrinsically mixing data and watermark through complex number arithmetic, the H[i][i]dden approach offers perfect reversibility and allows for embedding a watermark of in principle unlimited size, directly addressing key limitations of existing RDH techniques for individual and dispersed data. Two distinct protocols embodying this framework are proposed: the H[i][i]dden-EG protocol, providing joint RDH and encryption for individual data points by extending ElGamal encryption to Gaussian integers, and the H[i][i]dden-AggP protocol, which enables privacy-preserving data aggregation by taking advantage of a novel component-wise application of Paillier encryption to complex-valued data.
@@ Appendix A Modular Arithmetic over Gaussian Integers
It is worth pointing out that Gaussian prime numbers have some differences compared to the regular integer case. The obvious one is that Gaussian primes can have both real and imaginary parts. The second main difference is that not all integer primes are also Gaussian primes. For example, prime factorizations exists for both 2 and 5 in $\mathbb{Z}[i]$: 2=(1+i)(1i)2=(1+i)(1-i) and 5=(2+i)(2i)5=(2+i)(2-i). The integer primes $p$ that are also Gaussian primes satisfy the following condition $p\equiv 3\,(\operatorname{\mathrm{mod}}4)$ as proven, for example, in (Conrad, 2016).
@@
It is well known that, in regular integers, for a prime number $p\in\mathbb{Z}$, the elements of the cyclic group $\mathbb{Z}_{p}$ (or $\mathbb{Z}/p\mathbb{Z}$) are $0,1,\dots,p-1$, while the multiplicative cyclic group $\mathbb{Z}^{*}_{p}$ has the same elements, excluding 0. The order of the corresponding cyclic groups $\mathbb{Z}[i]_{\pi}$ and ${\mathbb{Z}[i]}^{*}_{\pi}$ is not so straightforward when $\pi$ is a Gaussian prime. For Gaussian integers, given a prime number $\pi\in\mathbb{Z}[i]$, the norm $\mathrm{N}(\pi)$ determines the order of the additive cyclic group $\mathbb{Z}[i]_{\pi}$, whereas $\mathrm{N}(\pi)-1$ is the order of the multiplicative cyclic group ${\mathbb{Z}[i]}^{*}_{\pi}$ (because of the exclusion of 0). Proofs for these facts can also be found in (Conrad, 2016).
As a practical example, the selection of representatives for the cyclic group ${\mathbb{Z}[i]}^{*}_{\pi}$ for $\pi=7$, which is both an integer and a Gaussian prime, is discussed below. First, we need to take into account that many possible selections of representatives exists, similar to the integer case. For example, 13 and 6 refer to the same equivalence class in both ${\mathbb{Z}}^{*}_{7}$ and ${\mathbb{Z}[i]}^{*}_{7}$, since 13 and 6 differ in a multiple of 7 (in fact, exactly 7). The only relevant consideration for the selection is not to include more than one representative for the same class. A method for such a selection is described in (Conrad, 2016) both when the divider has real and imaginary parts different from zero, and when the imaginary part is zero (the case of the Gaussian prime $\pi=7$).

### v3

@@ 1 Introduction
The ubiquitous generation and pervasive use of digital data stand as defining characteristics of the modern era. From vast networks of Internet of Things (IoT) devices and industrial sensors to intricate smart city infrastructures and critical healthcare systems, data fuels decision-making, automates processes, and enables unprecedented levels of connectivity. This reliance on data, often collected from distributed and potentially vulnerable environments, is critically linked to its trustworthiness (yan2014survey). Ensuring the confidentiality, integrity, authenticity, and privacy of information has become a priority, as compromised data can lead to erroneous conclusions, system failures, and severe breaches of privacy or security (liang2016review).
@@
Within this complex digital landscape, the concept of Reversible Data Hiding (RDH) (shi:2016; kumar2022reversible) emerges as a specialized and increasingly popular approach to data security. RDH involves embedding secret information (a “watermark”) directly into a host data item in such a way that, after extraction of the watermark, the original host data can be perfectly and exactly restored to its original state. This unique property, combining data embedding with guaranteed reversibility, makes RDH appropriate in sensitive applications such as medical imaging, legal documentation, and military intelligence, where any permanent alteration of the original data is unacceptable. Unlike traditional data encryption, which primarily ensures confidentiality, RDH focuses on providing intrinsic integrity, authenticity, and provenance, within the data itself.
@@
In the context of individual, scalar, or time-series numerical data, such as those originating from Wireless Sensor Networks (WSNs) (shi:2013), IoT devices (rahman:2021), or critical financial and medical monitoring systems (bahlearao:2019), RDH offers unique advantages. Foremost among these is its ability to ensure robust data integrity and authenticity at the granular level of single readings or sequential observations. By embedding a watermark that can be extracted and verified, receivers (e.g., a Data Collector in a WSN) can confirm that the data originated from a legitimate source and has not been tampered with in transit. This inherent capability to facilitate data provenance and traceability is particularly valuable, allowing for forensic analysis of data origin and processing paths in distributed and often resource-constrained environments.
@@
Beyond these general benefits, RDH offers distinct advantages compared to other commonly employed security primitives in certain contexts. Unlike cryptographic hashing or digital signatures, which typically append a separate authentication tag or signature to the data, RDH embeds the integrity verification mechanism directly within the data itself. This can simplify data handling and ensure the integrity proof travels jointly with the content. While conventional encryption ensures confidentiality, it does not directly provide data integrity; and Physical Unclonable Functions (PUFs) (gao2020physical) focus on device authentication rather than data integrity. RDH uniquely combines content-dependent integrity (through watermark verification) with the guarantee of perfect original data restoration, a feature not offered by one-way hash functions or irreversible authentication tags. This intrinsic embedding can be particularly beneficial in constrained environments or specialized data formats where appending additional metadata is undesirable or challenging.
@@
In a parallel development within cryptography, Homomorphic Encryption (HE) (acar2018survey) has emerged as a transformative technology for data privacy (iezzi2020practical). HE schemes allow computations to be performed directly on encrypted data without ever needing to decrypt it (zhang2020lvpda). The output of such a computation is an encrypted result that, when decrypted, is identical to the result of the same computation performed on the unencrypted data. This property enables the processing of sensitive information by untrusted third parties (such as cloud servers) while ensuring that the data’s confidentiality remains fully protected throughout its lifecycle.
@@ 1.1 Plan and Contributions of the Paper
  • A novel application of the extension of ElGamal encryption to the Gaussian integers within the H[i][i]dden-EG protocol, providing joint reversible information hiding and encryption specifically for individual data points.
@@ 2.1 Reversible Data Hiding
Reversible Data Hiding (shi:2016; kumar2022reversible) is a specialized field within information hiding that focuses on embedding secret information into host data in such a way that the original host data can be perfectly and exactly recovered after the hidden data (watermark) has been extracted. This property is essential in applications that do not allow the permanent alteration of the original content.
Historically, RDH techniques have been extensively developed and applied primarily to multimedia content, including digital images, audio, and video (shi:2016). Methods like Least Significant Bit (LSB) modification (when combined with original LSB storage) (celik:2002), Difference Expansion (DE) (tian:2003), Histogram Shifting (HS) (ni:2006), and Prediction-Error Expansion (PEE) (ou:2013) use the inherent statistical redundancy and the human perceptual system’s tolerance for slight alterations in these rich data formats. These approaches can embed information in an imperceptible mmanner while ensuring perfect reversibility, meaning the host data is faithfully restored to its original state after hidden data extraction.
@@
For this class of individual or dispersed data, existing reversible embedding methods tend to fall into a few categories. One common approach is Modulo or Remainder-based embedding, where the secret information $w$ (an integer in range $0,\ldots,q-1$) is embedded into a host integer $d$ using an operation like $d^{\prime}=q\cdot d+w$ (kabir2025rewap). A common special case involves embedding a single bit ($w\in\{0,1\}$) by setting $q=2$, resulting in $d^{\prime}=2d+w$ (kabir2024privacy). Upon extraction, $w$ is simply $d^{\prime}\,(\operatorname{\mathrm{mod}}q)$ and $d$ is $\lfloor d^{\prime}/q\rfloor$. When $q$ is a power of 2 ($q=2^{B}$), these extraction operations can be efficiently performed using bitwise logic (e.g., $w=d^{\prime}\text{ \& }(q-1)$ and $d=d^{\prime}\gg B$).
@@
Another relevant approach for sequential data is Difference Expansion (DE), or its more general form, Prediction-Error Expansion (shi:2013). These methods exploit relationships between consecutive data samples by computing prediction errors (e.g., $e_{k}=d_{k}-d_{k-1}$) and then embedding information by expanding these errors.
@@ 2.2 Homomorphic Encryption
Homomorphic Encryption (acar2018survey) is a unique and powerful class of cryptographic schemes that allows a user to perform computations on encrypted data without first having to decrypt it. The result of these operations is an encrypted output which, when decrypted, is the same as if the operations had been performed on the original, unencrypted data. This remarkable capability is useful for maintaining data privacy and confidentiality when computations are outsourced to untrusted third parties, such as cloud services or remote data aggregators.
HE schemes are typically classified based on the types and number of operations they support. Partially Homomorphic Encryption (PHE) schemes, such as (cominetti2020fast), support an unlimited number of a single type of operation, either addition or multiplication. Somewhat Homomorphic Encryption (SHE) schemes, which include the pioneering work by Boneh-Goh-Nissim (BGN) in (boneh2005evaluating), can perform a limited number of both additions and multiplications. Finally, Fully Homomorphic Encryption (FHE) schemes, such as Ring Learning with Errors (RLWE) (brakerski2011fully), support an unlimited number of both operations, allowing for arbitrary computations on encrypted data. In this paper, two classical PHE schemes, elgamal1985public and paillier1999public, are applied for their distinct homomorphic properties that enable the proposed protocols for both individual data encryption and privacy-preserving data aggregation.
@@ 2.2.1 ElGamal Encryption
ElGamal encryption (elgamal1985public) is an asymmetric key encryption algorithm based on the Diffie-Hellman key exchange scheme. Its security relies on the difficulty of the discrete logarithm problem (DLP) in the multiplicative group of integers modulo a prime number.
@@ 2.2.2 Paillier Encryption
paillier1999public is a probabilistic asymmetric encryption algorithm. Unlike ElGamal, which is homomorphic with respect to multiplication, Paillier encryption is additively homomorphic, meaning that sums of plaintexts can be computed from their ciphertexts.
@@ 3.1 Basic Definitions
Unlike reversible data hiding for multimedia content or highly redundant data —where source redundancy allows for techniques such as difference expansion, histogram shifting, or prediction error expansion to hide information without necessarily increasing carrier size— embedding a watermark $w_{k}$ in a single scalar $d_{j,k}$ such that both can be perfectly recovered requires data expansion. For instance, as seen in kabir2025rewap, embedding a $B$-bit watermark using an operation like $d^{\prime}_{j,k}=2^{B}\cdot d_{j,k}+w_{k}$ increases the bit-length of the original data by exactly $B$ bits. From an information-theoretic perspective, if the host data consists of $A$ bits and the watermark comprises $B$ bits, and both are independent and occupy their full possible ranges, the combined state space consists of $2^{A+B}$ distinct values. In such cases, the result must be encoded with at least $A+B$ bits for perfect recovery. While “slack” in a data’s restricted range (e.g., values $d_{j,k}\in[101,200]$ encoded in a 1-byte container) can theoretically host bits without expansion, scalar reversible embedding of independent, full-range variables fundamentally requires increasing the bit-length of the transmitted values. $\triangle$
@@ 3.2 Proposed Reversible Watermarking Scheme
The proposed method is vulnerable against Gaussian integer factorization. Even with just one eavesdropped value, a malicious observer may try to compute the Gaussian integer factors of the watermarked data and obtain knowledge about the original data, the watermark or the watermarking key. Consequently, for practical applications, it is not recommended to rely solely on the challenge $\lambda_{k}$ as a security measure. The scheme can be combined with random numbers and/or cryptographic functions for increased security and privacy, as shown in the protocols proposed in Section 5. $\triangle$
@@
Hence, an eavesdropper who accessed $\delta^{\prime}_{j,k}$ would not be able to determine the data or the watermark directly, but may try to infer information from Gaussian integer factorization, as discussed in Remark 8. In addition, if the eavesdropper had access to the watermarked values corresponding to different time samples ($k$) or different devices ($j$), they could try to obtain even more information about $\lambda_{k}$, especially if it is not changed frequently. $\triangle$
@@ 4 Homomorphic Encryption on the Complex Domain
The extension of ElGamal encryption to Gaussian integers, first presented in El-Kassar2001El-Gamal, requires a careful examination of the mathematical basis of the cryptosystem. The first building block of the standard ElGamal cryptosystem is the multiplicative cyclic group $\mathbb{Z}^{*}_{p}$, for a prime number $p$, and a generator $g$ for such a group. In the Gaussian integer case, we must have a prime number $\pi$ in $\mathbb{Z}[i]$ and a consistent definition of the modular arithmetic $(\operatorname{\mathrm{mod}}\pi)$.
For a better understanding of the adaptation of this cryptosystem to the complex domain, a brief introduction to modular arithmetic over Gaussian integers is provided in A. A more detailed and comprehensive reference on Gaussian integers can be found in (conrad2016gaussian).
@@ 4.1 ElGamal Encryption over Gaussian Integers
As shown in El-Kassar2001El-Gamal, the principles of ElGamal encryption can be extended to other finite cyclic groups, including the multiplicative group of Gaussian integers modulo a Gaussian prime. This adaptation exploits the rich algebraic structure of Gaussian integers while maintaining the security reliance on the DLP.
@@
As a key property, the order of this group, $\mathcal{N}=p^{2}-1$, is significantly larger than p1p-1 for the same numerical prime $p$ used in the regular integer ElGamal. This relationship, $\mathcal{N}=(p-1)(p+1)$, is particularly advantageous for security. If a prime $p$ is chosen such that p1p-1 possesses a large prime factor (a standard requirement for robust security in integer-based ElGamal against algorithms like Pohlig-Hellman (pohlig:1978)), then this same large prime factor will inherently be a factor of $\mathcal{N}$. Therefore, the Gaussian integer ElGamal variant naturally inherits at least the same level of security (in terms of resistance to Pohlig-Hellman attacks) as its regular integer counterpart for a given $p$. Furthermore, the existence of the additional factor $(p+1)$ in $\mathcal{N}$ can potentially introduce even larger prime factors, further enhancing the security of the scheme. Thus, careful selection of the Gaussian prime $p$ to ensure that the largest prime factor of $\mathcal{N}$ is sufficiently large guides robust parameter choices.
@@ 7 Conclusion
This paper introduces Hiding in the Imaginary Domain with Data Encryption (H[i][i]dden), a novel framework for RDH that operates uniquely within the complex domain. By intrinsically mixing data and watermark through complex number arithmetic, the H[i][i]dden approach offers perfect reversibility and allows for embedding a watermark of in principle unlimited size, directly addressing key limitations of existing RDH techniques for individual and dispersed data. Two distinct protocols embodying this framework are proposed: the H[i][i]dden-EG protocol, providing joint RDH and encryption for individual data points by using the extension of ElGamal encryption to Gaussian integers, and the H[i][i]dden-AggP protocol, which enables privacy-preserving data aggregation by taking advantage of a novel component-wise application of Paillier encryption to complex-valued data.
@@ Appendix A Modular Arithmetic over Gaussian Integers
It is worth pointing out that Gaussian prime numbers have some differences compared to the regular integer case. The obvious one is that Gaussian primes can have both real and imaginary parts. The second main difference is that not all integer primes are also Gaussian primes. For example, prime factorizations exists for both 2 and 5 in $\mathbb{Z}[i]$: 2=(1+i)(1i)2=(1+i)(1-i) and 5=(2+i)(2i)5=(2+i)(2-i). The integer primes $p$ that are also Gaussian primes satisfy the following condition $p\equiv 3\,(\operatorname{\mathrm{mod}}4)$ as proven, for example, in (conrad2016gaussian).
@@
It is well known that, in regular integers, for a prime number $p\in\mathbb{Z}$, the elements of the cyclic group $\mathbb{Z}_{p}$ (or $\mathbb{Z}/p\mathbb{Z}$) are $0,1,\dots,p-1$, while the multiplicative cyclic group $\mathbb{Z}^{*}_{p}$ has the same elements, excluding 0. The order of the corresponding cyclic groups $\mathbb{Z}[i]_{\pi}$ and ${\mathbb{Z}[i]}^{*}_{\pi}$ is not so straightforward when $\pi$ is a Gaussian prime. For Gaussian integers, given a prime number $\pi\in\mathbb{Z}[i]$, the norm $\mathrm{N}(\pi)$ determines the order of the additive cyclic group $\mathbb{Z}[i]_{\pi}$, whereas $\mathrm{N}(\pi)-1$ is the order of the multiplicative cyclic group ${\mathbb{Z}[i]}^{*}_{\pi}$ (because of the exclusion of 0). Proofs for these facts can also be found in (conrad2016gaussian).
As a practical example, the selection of representatives for the cyclic group ${\mathbb{Z}[i]}^{*}_{\pi}$ for $\pi=7$, which is both an integer and a Gaussian prime, is discussed below. First, we need to take into account that many possible selections of representatives exists, similar to the integer case. For example, 13 and 6 refer to the same equivalence class in both ${\mathbb{Z}}^{*}_{7}$ and ${\mathbb{Z}[i]}^{*}_{7}$, since 13 and 6 differ in a multiple of 7 (in fact, exactly 7). The only relevant consideration for the selection is not to include more than one representative for the same class. A method for such a selection is described in (conrad2016gaussian) both when the divider has real and imaginary parts different from zero, and when the imaginary part is zero (the case of the Gaussian prime $\pi=7$).